### v1

# Distributed Thompson sampling under constrained communication

Abstract
In Bayesian optimization, a black-box function is maximized via the use of a surrogate model. We apply distributed Thompson sampling, using a Gaussian process as a surrogate model, to approach the multi-agent Bayesian optimization problem. In our distributed Thompson sampling implementation, each agent receives sampled points from neighbors, where the communication network is encoded in a graph; each agent utilizes a Gaussian process to model the objective function. We demonstrate a theoretical bound on Bayesian Simple Regret, where the bound depends on the size of the largest complete subgraph of the communication graph. Unlike in batch Bayesian optimization, this bound is applicable in cases where the communication graph amongst agents is constrained. When compared to sequential Thompson sampling, our bound guarantees faster convergence with respect to time as long as there is a fully connected subgraph of at least two agents. We confirm the efficacy of our algorithm with numerical simulations on traditional optimization test functions, illustrating the significance of graph connectivity on improving regret convergence.
A Introduction
Bayesian optimization (BO) [1, 2] is a technique used to optimize a black-box objective function, where the function is unknown and expensive to evaluate. BO is particularly useful in applications where sampling the objective function involves significant computational resources, such as robotics [3], [4], hyperparameter tuning in machine learning [5], [6], and simulation-based optimization [7]. In BO, a surrogate model is used to approximate the unknown objective function - a Gaussian process (GP) [8] is commonly used for this task. An acquisition function based on this surrogate model is used to determine which points should be queried next. Using the acquisition function, query points can be chosen to incorporate a tradeoff between exploration and exploitation by balancing uncertainty reduction in the model and sampling near the current optimum.
Acquisition functions govern the query selection process, which is critical for ensuring efficient optimization. A popular acquisition function is expected improvement (EI), within which the next query point is selected as one which maximizes the expected improvement over the current best-known value [2]. Alternatively, an upper confidence bound (UCB) can be used to select the next query point by selecting a point which maximizes an upper bound on the predicted value [9]. In entropy search (ES), the query points are selected such that the mutual information between the query points and the location of the global optimum is maximized [10], [11]. In Thompson sampling, the next query point is selected as one which maximizes a realization sampled from the posterior of the surrogate model [12]. Thompson sampling is of particular interest to us, as it lends itself well to multi-agent settings [13].
Evaluation of the efficacy of an implementation of Bayesian optimization is multifaceted. Standard performance metrics for Bayesian optimization include time complexity, sample complexity, and regret. Time complexity describes the number of iterations the algorithm requires to converge to a sufficiently optimal solution [14]. Sample complexity refers to the quantity of samples required for the algorithm to reach an approximately optimal solution [15]. To evaluate the quality of the selection of query points, regret is often used, which quantifies the gap between sampled points and the global optimum [16]. Regret can take many forms, including simple regret, which considers the performance of the best queried point [17], and cumulative regret, which considers performance of all prior sampled points [18].
Bayesian optimization can be used in a multi-agent setting, in which multiple agents can sample the objective function at a single timestep. This is advantageous because in multi-agent settings, the extrema of the objective function can be determined in fewer timesteps than in a single-agent setting - however, additional complexity is introduced, because in order to fully realize this advantage, the agents must engage in collaboration. To facilitate collaboration, existing literature has been focused on batch Bayesian optimization, a centralized setting in which a central coordinator has access to each agents’ acquired information [19], [13]. It then processes and computes the sampling decisions for all agents, which are in turn communicated to each agent. These decisions are often disseminated in batches, allowing multiple agents to simultaneously sample points and consequently parallelizing the optimization process [20], [21].
Centralized approaches are inapplicable in distributed cases, in which there is no centralized coordinator and each agent must possess a local instance of the optimization algorithm [22]. Additionally, they often do not scale well, as they require a central coordinator to manage the processing of all agents’ data. Distributed networks are prevalent in real-world applications, such as in multi-robot source seeking and sensor networks [19], [23]. It may not be the case that all agents have access to all prior sampled points as in the batch setting - communication may be constrained, where some agents are only able to communicate with specific other agents [24]. These constraints may be due to limited communication capacity or computational capacity of the agents, or due to physical proximity constraints. In this work, we study the distributed multi-agent case, where the communication network of agents is given by a static graph.
Our contribution: we propose a distributed Thompson sampling algorithm to approach the multi-agent Bayesian optimization problem. Unlike in batch cases, in which all agents query points based on a posterior distribution which depends on all previously sampled points, our algorithm is applicable in cases with constrained communication. Prior literature providing theoretical guarantees on distributed Thompson sampling algorithms require fully connected communication graphs, even in asynchronous cases [13], [25], and thus are inapplicable in settings with constrained communication. We establish a theoretical bound on regret that is dependent on the size of the largest complete subgraph of the communication graph - we are the first to demonstrate a distributed BO algorithm with probable guarantees in the constrained communication case. Additionally, we establish the efficacy of our performance and confirm the feasibility of our theoretical bound with numerical simulations on standard optimization test functions.
B Problem Formulation and Preliminaries
B-A Problem Formulation
For a compact set $\mathcal{X}\subset\mathbb{R}^{d}$, consider an unknown continuous function $f:\mathcal{X}\to\mathbb{R},$ with optimizer $x^{*}$. In the BO setting, we wish to find the maximum of this function, where we are only able to sample $f$ through expensive and noisy evaluations. We assume any of $M$ agents can query $f$ at any point and receive a noisy value $y=f(x)+\epsilon,$ with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}).$ Agents query $f$ throughout a total of $T$ iterations. For agent $i\in\{1,\ldots,M\}$ and iteration $t\in\{1,\ldots,T\},$ $x_{t,i}$ is the query point, and $y_{t,i}$ is the corresponding evaluation. Define $X_{t,i}=\{x_{1,i},\ldots,x_{t,i}\},$ $Y_{t,i}=\{y_{1,i},\ldots,y_{t,i}\}$ to be the queries and evaluations made by agent $i$ up to time $t.$ The communication network of $M$ agents is described by graph $G=(V,E),$ where $|V|=M,$ and $E\subset\{\{i,j\}:i,j\in V,i\neq j\}.$ An unordered pair $\{i,j\}\in E$ if agents $i$ and $j$ are able to communicate with each other. Additionally, we denote the set of neighbors of agent $i$ as $N(i)=\{j:\{i,j\}\in E\}.$ The data accessible to agent $i$ at time $t$ is $D_{t,i}=\{(x_{\tau,j},y_{\tau,j})\}_{j\in N(i)\cup i,\tau<t}.$ The set $D_{t,i}$ contains all sampled points by agent $i$ and its neighbors.
B-B Gaussian Process
We will use a Gaussian process (GP) to model our unknown objective function $f$ in our BO setting. Recall the unknown continuous objective function $f:\mathcal{X}\to\mathbb{R}.$ Let $\mathbf{X}_{D_{t}}=\{x_{1},x_{2},\cdots,x_{t}\},$ where $x_{j}$ is the $j$th evaluated point. Let $k:\mathcal{X}^{2}\to\mathbb{R}$ be a kernel function selected based on prior beliefs about $f$. Define
$\mu_{D_{t}}(x)$ $=\mathbf{k}_{t}(x)^{\intercal}(\mathbf{K}_{D_{t}}+\sigma^{2}_{n}\mathbf{I})^{-1}\mathbf{y}_{D_{t}}$
$k_{D_{t}}(x,x^{\prime})$ $=k(x,x^{\prime})-\mathbf{k}_{D_{t}}(x)^{\intercal}(\mathbf{K}_{D_{t}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{D_{t}}(x^{\prime}),$
where $\mathbf{K}_{D_{t}}:=[k(x^{\prime},x^{\prime\prime})]_{x^{\prime},x^{\prime\prime}\in\mathbf{X}_{D_{t}}},$ $\mathbf{k}_{D_{t}}(x):=[k(x^{\prime},x)]_{x^{\prime}\in\mathbf{X}_{D_{t}}}$ and $\mathbf{y}_{D_{t}}=\{f(x^{\prime})+\epsilon^{\prime}\}_{x^{\prime}\in\mathbf{X}_{D_{t}}},$ where $\epsilon^{\prime}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}).$ Thus we can define our GP, in which we denote $f|\mathcal{F}_{D_{t}}\sim GP(\mu_{D_{t}}(x),k_{D_{t}}(x,x^{\prime})).$ Note that due to the nature of the $GP$, it is the case that for any $x\in\mathcal{X},$ $f(x)|\mathcal{F}_{D_{t}}\sim N(\mu_{D_{t}}(x),\sigma_{D_{t}}^{2}(x)),$ where $\sigma_{D_{t}}^{2}(x)=k_{D_{t}}(x,x)$ [8]. Furthermore, recall the distributed multi-agent setting, where each of $M$ agents have access to queried points in set $D_{t,i},$ where $D_{t,i}$ and $D_{t,j}$ may not be equal for distinct agents $i$ and $j.$ In our distributed setting, we equip each agent with a unique GP model of $f$, and thus we write $f_{i}(x)|\mathcal{F}_{t,i}\sim\mathcal{N}(\mu_{D_{t,i}}(x),\sigma_{D_{t,i}}^{2}(x))$
B-C Regret
Our metric for algorithm performance is regret, which is an assessment of the quality of sampled points. Specifically, we consider the simple regret, which is the difference between the optimal value of the true function and the best value achieved amongst the previous queried points. This definition of regret is useful because optimization settings focus on locating the extrema of a function, and the simple regret tracks the smallest gap between the value at a sampled point and the optimal value. To account for randomness of $f$ in our regret expression, we take expectation of simple regret to yield the following expression, which is called Bayesian Simple Regret:
$R_{SB}(t)=\min_{i\in\{1,2,\cdots,M\},\tau\in\{1,2,\cdots,t\}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$ (1)
In our theoretical analysis, we provide a bound on Bayesian Simple Regret.
B-D Thompson Sampling
Thompson sampling is an algorithm for sequential decision making that can be utilized in this context for determining the next point of the objective function to query [12]. When using Thompson sampling in our Bayesian optimization framework, a function is sampled from the posterior distribution of the Gaussian process. The maximizer of this function is the next query point at which the black-box objective function is sampled. The Gaussian process is then updated with new information from this sample, and the process repeats for the duration of the experiment.
In sequential Thompson sampling, each subsequent query point is determined based on a model updated on all prior sampled points. Alternative implementations of Thompson sampling include batch Thompson sampling, in which multiple query points are determined as a set at each round, and query the objective function in parallel [13], [21]. Batch Thompson sampling is advantageous in systems capable of parallelizing, e.g. multi-agent systems, because it allows for convergence in fewer number of rounds than sequential Thompson sampling.
In the context of multi-agent systems, Batch Thompson sampling can be thought of as a centralized network, where all agents are simultaneously sampling the objective function at different query points, and all agents have access to the same information. However, this may not be realistic in real-world situations, where communication between agents may be more constrained due to bandwidth limitations, computational constrictions, or privacy concerns. In these cases, agents may only have access to the sampled points by some subset of the other agents, and thus each agent will have may have different models of the objective function across their unique GPs. We propose a Thompson Sampling algorithm for this constrained communication case, and provide theoretical guarantees for the algorithm.
C Algorithm: Distributed Thompson Sampling
In our implementation of distributed Thompson sampling, each of $M$ agents have distinct Gaussian processes $\mathcal{GP}_{i}$ for modeling the objective function. At each time step $t,$ all agents update their GPs with the data history available to them. The agent then queries the objective function at $x_{t,i},$ which is the maximizer of the acquisition function sampled from the posterior GP, $\hat{f}_{t,i}\sim\mathcal{GP}_{t,i}.$ Each agent then communicates its sampled point to its neighbors, receives the points sampled by their neighbors, and updates their data history accordingly. The collection of data received by neighbors of agent $i$ at time $t$ is denoted as $C_{t,i}=\{(x_{t,j},y_{t,j})\}_{j\in N(i)}.$ Our method is shown in Algorithm 1.
Algorithm 1 Distributed Thompson Sampling
1:Place GP prior on $f$
2:for i$=1,\ldots,M$ do
3:     Initial data $D_{1,i}$
4:     $\mathcal{GP}_{0,i}\leftarrow GP$
5:end for
6:for $t=1,\ldots,T$ do
7:     for i$=1,\ldots,M$ do
8:         Update posterior $\mathcal{GP}_{t,i}$ conditioned on $D_{t,i}$
9:         Sample $\hat{f}_{t,i}\sim\mathcal{GP}_{t,i}$
10:         Choose next query point
11:                  $x_{t,i}\leftarrow\arg\max_{x}\hat{f}_{t,i}(x)$
12:         Observe $y_{t,i}$
13:         Broadcast $y_{t,i}$ to neighbors $N(i)$;
14:         Collect evaluations $C_{t,i}$ from neighbors $N(i)$
15:         Update data history $D_{t+1,i}\leftarrow D_{t,i}\cup C_{t,i}\cup\{(x_{t,i},y_{t,i})\}$
16:     end for
17:end for
C-A Theoretical Result
We analyze the performance of the Distributed Thompson Sampling algorithm on the Bayesian Simple Regret metric. Our regret bound depends on the number of timesteps $T$ and the structure of the agent communication graph G. As in prior work, we utilize notions from information theory in our regret bound [26].
Our regret bound involves the Maximum Information Gain (MIG), which is a constant that captures the difficulty of the objective function. MIG has been shown to be bounded for several kernel functions commonly used with GPs, including Squared Exponential and Matérn kernels, the latter of which was used in our numerical implementation [9].
Let $D=\{x_{1},\ldots,x_{t}\}\subset\mathcal{X},$ and define $y_{D}=\{(x,f(x)+\epsilon):x\in D\}.$ The MIG is denoted as
$\Psi_{t}=\max_{D\subset\mathcal{X},|D|=t}I(f;y_{D}),$ (2)
where $I$ is the Shannon Mutual Information. The MIG $\Psi_{t}$ represents the largest mutual information gain from $f$ by sampling $t$ points. Additionally, for any positive integer $n$, we define the constant $\xi_{n},$ which bounds the information gain of the current round of evaluations [13]. Suppose $|D|=t$ points were already sampled, and $i$ points are being queried in the current round of evaluations, with $i<n;$ denote these points in set $A,$ where $A\subset\mathcal{X},$ and $y_{A}=\{(x,f(x)+\epsilon):x\in A\}.$ Then for $i\geq 1,$ $\xi_{n}$ satisfies
$\max_{A\subset\mathcal{X},|A|<n}I(f;y_{A}|y_{D})\leq\frac{1}{2}\log(\xi_{n}).$ (3)
In the following theorem, we provide a bound for Bayesian Simple Regret for an $M$ agent system with communication graph $G.$ After proving Theorem 3.1, we will introduce and discuss its implication, Corollary 3.2, which is our main convergence result.
Theorem 3.1.
Suppose $k(x,x^{\prime})\leq 1$ for all $x,x^{\prime}.$ Let $G_{s}=(V_{s},E_{s})$ be a complete subgraph of $G.$ Then the Bayesian Simple Regret after $t$ timesteps satisfies $R_{SB}(t)\leq\frac{C_{1}}{t|V_{s}|}+\sqrt{\frac{C_{2}\xi_{|V_{s}|}\beta_{t}\Psi_{t|V_{s}|}}{t|V_{s}|}},$ where $\beta_{t}=2\log(t^{2}|V_{s}||\mathcal{X}|),$ $C_{1}=\frac{\sqrt{2}\pi^{3/2}}{12},$ and $C_{2}=\frac{2}{\log(1+\sigma_{\epsilon}^{-2})}.$
Proof. The structure of our proof follows techniques from Kandasamy et al. [13]. We aim to provide a bound on Bayesian Simple Regret. Our proof begins by noting that we can bound Bayesian Simple Regret the average expected regret of any subset of agents. We then decompose this into three sums, each of which utilize a confidence function $U_{t,i}(\cdot).$ We bound each of these sums by incorporating notions from information theory, which allows us to use information gain constants introduced in Equations 3 and 2 to analyze the efficacy of the Gaussian process.
Recall from equation 1 that $R_{SB}(t)=\min_{i\in\{1,2,\cdots,M\},\tau\in\{1,2,\cdots,t\}}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})].$ Define $U_{t,i}(x)=\mu_{D_{t,i}}(x)+\beta_{t}^{1/2}\sigma_{D_{t,i}}(x).$ Note that, since $R_{SB}(t)$ represents a minimum value of the expression $\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})],$ it is upper bounded by the average of this expression across any subset of agents. Enumerate the agents in the complete graph $G_{s}=(V_{s},E_{s})$ as $\{1,2,\cdots,|V_{s}|\},$ and thus we reindex the $M$ agents such that agents $i\in\{1,2,\cdots,|V_{s}|\}\in V_{s}$ and agents $j\in\{|V_{s}|+1,\cdots,M\}\in V\backslash V_{s}.$ Thus we can write that
$R_{SB}(t)\leq\frac{1}{t|V_{s}|}\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})].$
We will then focus on upper bounding the expression $\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})].$
$\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})]$
$=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-U_{\tau,i}(x^{*})+U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})$
$\hskip 56.9055pt+U_{\tau,i}(x_{\tau,i})-f_{i}(x_{\tau,i})]$
$=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\underbrace{\mathbb{E}[f_{i}(x^{*})-U_{\tau,i}(x^{*})]}_{S1}+\underbrace{\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})]}_{S2}$
$\hskip 56.9055pt+\underbrace{\mathbb{E}[U_{\tau,i}(x_{\tau,i})-f_{i}(x_{\tau,i})]}_{S3}$
We will now bound each of these sums.
S1. Let’s begin by upper bounding the sum                       $S1=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-U_{\tau,i}(x^{*})].$
$S1$ $=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-U_{\tau,i}(x^{*})]$ (4)
$\leq\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\mathbb{I}\{f_{i}(x^{*})>U_{\tau,i}(x^{*})\}\left(f_{i}(x^{*})-U_{\tau,i}(x^{*})\right)]$ (5)
$\leq\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sum_{x\in\mathcal{X}}\frac{e^{-\beta_{\tau}/2}}{\sqrt{2\pi}}$ (6)
$\leq\sum_{\tau=1}^{t}\frac{1}{\tau^{2}\sqrt{2\pi}}\leq\frac{\sqrt{2}\pi^{3/2}}{12}$ (7)
Line (5) upper bounds $S1$ by positive terms, Line (6) utilizes Lemma A.1 in Appendix A to bound the expectation of positive terms in a normal distribution. Line (7) results from substituting for $\beta_{\tau},$ and the fact that $\sum_{j=1}^{\infty}\frac{1}{j^{2}}=\frac{\pi^{2}}{6}.$ Therefore, we have established an upper bound for $S1$.
S2. Let’s evaluate the expression
$S2=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})].$ First let’s focus on the interior of the summation, $\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})].$ We will proceed to show that this expression evaluates to $0$
By the law of total expectation, $\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})]=\mathbb{E}[\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})]|D_{t,i}]$. Because $x_{t,i}$ is sampled from the posterior distribution of $x^{*}|D_{t,i},$ we must have that $x_{t,i}$ and $x^{*}$ have the same distribution after conditioning on the acquired data, and thus $x_{t,i}|D_{t,i}\sim x^{*}|D_{t,i}.$ We also notice that $U_{t,i}$ is deterministic when conditioned on $D_{t,i}.$ Thus we have that $\mathbb{E}[\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})]|D_{t,i}]]=0,$ and consequently $S2=0.$
S3. Lastly, we bound
$S3=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[U_{\tau,i}(x_{\tau,i})\!-\!f_{i}(x_{\tau,i})].$ For the evaluation of this sum, we will introduce some additional notation. Denote $\bar{D}_{t}=\{(x_{\tau,j},y_{\tau,j})\}_{j\in V_{s},\tau<t}.$ We can think of $\bar{D}_{t}$ as representing the data acquired by round $t$ in a batch setting exclusively by agents in $V_{s}$. Additionally, define $\sigma_{t,i}(x):=\sigma(x)\mid\bar{D}_{t}\cup\left\{(x_{t,j},y_{t,j})\right\}_{j\in V_{s},j<i}.$ We pick to define $\sigma_{t,i}(\cdot)$ in this way to impose an ordering of the acquired data; in this sense, $\sigma_{t,i}(\cdot)$ depends on $(t-1)|V_{s}|+i-1$ previously sampled points. With this additional notation, we are well-equipped to upper bound $S3.$
$S3$ $=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[U_{\tau,i}(x_{\tau,i})-f_{i}(x_{\tau,i})]$ (8)
$=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\mathbb{E}[\mu_{D_{\tau,i}}(x_{\tau,i})+\beta_{\tau}^{1/2}\sigma_{D_{\tau,i}}(x_{\tau,i})$
$\hskip 49.79231pt-f_{i}(x_{\tau,i})]|D_{\tau,i}]]$ (9)
$=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\beta_{\tau}^{1/2}\sigma_{D_{\tau,i}}(x_{\tau,i})]$ (10)
$\leq\beta_{t}^{1/2}\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\sigma_{D_{\tau,i}}(x_{\tau,i})]$ (11)
$\leq\beta_{t}^{1/2}\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\sigma_{\bar{D}_{\tau}}(x_{\tau,i})]$ (12)
$\leq\beta_{t}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sigma_{\tau,i}(x_{\tau,i})\exp(I(f;\{y_{\tau,j}\}_{j<i}|y_{\bar{D}_{\tau}})\right]$ (13)
$\leq\beta_{t}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sigma_{\tau,i}(x_{\tau,i})\xi_{i}^{1/2}\right]$ (14)
$\leq\beta_{t}^{1/2}\xi_{|V_{s}|}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sigma_{\tau,i}(x_{\tau,i})\right]$ (15)
$\leq\beta_{t}^{1/2}\xi_{|V_{s}|}^{1/2}\mathbb{E}\left[\Big{(}t|V_{s}|\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sigma^{2}_{\tau,i}(x_{\tau,i})\Big{)}^{1/2}\right]$ (16)
$\leq\beta_{t}^{1/2}\xi_{|V_{s}|}^{1/2}\sqrt{\frac{2t|V_{s}|\Psi_{t|V_{s}|}}{\log(1+\sigma_{\epsilon}^{-2})}}$ (17)
$\leq\sqrt{\frac{2\xi_{|V_{s}|}t|V_{s}|\beta_{t}\Psi_{t|V_{s}|}}{\log(1+\sigma_{\epsilon}^{-2})}}$ (18)
Lines (10) and (11) follow by the law of total expectation. Line (12) follows by noting that $\beta_{\tau}$ is increasing with $\tau.$ Recall that because $G_{s}$ is a subgraph of $G,$ for all agents $i\in V_{s},$ $V_{s}\subset(N(i)\cup i).$ Therefore, $\bar{D}_{\tau}$ is contained in $D_{\tau,i}$ for all agents $i\in V_{s}.$ In essence, $\bar{D}_{\tau}\subset D_{\tau,i}\implies\sigma_{\bar{D}_{\tau}}\geq\sigma_{D_{\tau,i}}.$ Line (13) applies this property. Line (14) follows from Lemma A.4 in Appendix A, which introduces $\xi_{i}$ to bound the information gain of the current set of evaluations. Line (15) follows from the definition of $\xi_{i},$ which was stated in Equation 3. Line (16) is a consequence of the fact that $\xi_{i}$ is increasing with $i,$ and Line (17) follows from application of the Cauchy-Schwarz inequality. Finally, Line (18) is a consequence of Lemma A.3, which bounds the sums of the posterior variances by the MIG term.
Figure 1: Simulations of regret analysis with varying agent quantity on star graph configuration, on Ackley (left) and Rosenbrock (right) objective functions.
Recall that $\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})]=S1+S2+S3.$ Thus, we have shown that $\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})]\leq\frac{\sqrt{2}\pi^{3/2}}{12}+\sqrt{\frac{2\xi_{|V_{s}|}t|V_{s}|\beta_{t}\Psi_{t|V_{s}|}}{\log(1+\sigma_{\epsilon}^{-2})}}.$ However, we wish to bound $R_{SB}(t).$ Since $R_{SB}(t)\leq\frac{1}{t|V_{s}|}\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})],$ it directly follows that $R_{SB}(t)\leq\frac{\sqrt{2}\pi^{3/2}}{12t|V_{s}|}+\sqrt{\frac{2\xi_{|V_{s}|}\beta_{t}\Psi_{t|V_{s}|}}{t|V_{s}|\log(1+\sigma_{\epsilon}^{-2})}},$ thus concluding our proof.
Picking $G_{s}$ to be the largest complete subgraph of the communication network $G$ yields the following corollary.
Corollary 3.2.
Suppose $k(x,x^{\prime})\leq 1$ for all $x,x^{\prime}.$ Let $G_{max}=(V_{max},E_{max})$ be the largest complete subgraph of $G.$ Then the Bayesian Simple Regret after $t$ timesteps satisfies $R_{SB}(t)\leq\frac{C_{1}}{t|V_{max}|}+\sqrt{\frac{C_{2}\xi_{|V_{max}|}\beta_{t}\Psi_{t|V_{max}|}}{t|V_{max}|}},$ where $\beta_{t}=2\log(t^{2}|V_{s}||\mathcal{X}|),$ $C_{1}=\frac{\sqrt{2}\pi^{3/2}}{12},$ and $C_{2}=\frac{2}{\log(1+\sigma_{\epsilon}^{-2})}.$
From the corollary, we see that $R_{SB}(t)=\tilde{O}\left(\sqrt{\frac{\xi_{|V_{max}|}\Psi_{t|V_{max}|}}{t|V_{max}|}}\right)$. We note that the term $\Psi_{t|V_{max}|}$ corresponds to the maximal mutual information gain from $t|V_{max}|$ observations, and that this quantity depends only logarithmically on $t|V_{max}|$ for standard kernels such as the squared exponential kernel. For more details, we refer to Appendix B. We note that the term $\xi_{|V_{max}|}$ is the price we pay for the delay in communication within the subgraph $G_{max}$, and is a standard term that arises in non-sequential Bayesian Optimization. By an appropriately designed initialization phase, this term can actually be reduced to $\tilde{O}(1)$ (see Appendix B.3 in [21] for more details). Thus, compared to the sequential case with $t$ rounds which has simple regret $\tilde{O}\left(\sqrt{\frac{1}{t}}\right)$ [9], our algorithm satisfies a regret of $\tilde{O}\left(\sqrt{\frac{1}{t|V_{max}|}}\right)$, i.e. an improvement of $\sqrt{\frac{1}{|V_{max}|}}$, demonstrating the benefit of the network structure.
D Numerical Experiments
D-A Simulation
[…]
$R_{B}(t)=f^{*}-\max_{\begin{subarray}{c}i\in\{1,2,\ldots M\}\\ {{x}_{t,i}\in X_{t,i}}\end{subarray}}f(\mathbf{x}_{t}^{i}),\hskip 7.11317pt\overline{R_{B}}(t)=\sum_{\tau=1}^{t}R_{B}(\tau),$
where $f^{*}=\max_{x\in\mathcal{X}}f(x).$sayInstant worst regret $R_{W}$ is the regret of the worst performing agent at time step $t,$ and we define its sum to be the quantity $\overline{R_{W}}:$
$R_{W}(t)=f^{*}-\min_{i\in\{1,2,\ldots M\}}f(x_{t,i}),\hskip 7.11317pt\overline{R_{W}}(t)=\sum_{\tau=1}^{t}R_{W}(\tau).$
[…]
Instant average regret $R_{A}$ is the average of the regrets of all agents at time step $t,$ and we define its sum to be $\overline{R_{A}}:$
[…]
D-B Discussion
Our Distributed Thompson Sampling algorithm was able to achieve the extrema of the Ackley and Rosenbrock objective functions, and thus is effective at the Bayesian optimization task. Our theoretical result bounds expectation of Bayesian Simple Regret, $R_{SB}(t),$ with the bound dependent on the size of the largest complete subgraph of the communication network between agents. In the star graph configuration, the largest complete subgraph contained $2$ vertices for any star graph containing at least $2$ vertices. Interestingly, in Figure 1, we see similar performance in the $3,5,$ and $7$ agent cases, with significantly worse performance in the single agent case. This matches intuition motivated by our theoretical result, as the $3,5,$ and $7$ agent cases all have largest complete subgraph size of $2,$ as opposed to the single agent case, where the largest complete subgraph is of size $1$
E Conclusion
In this paper, we proposed a Distributed Thompson Sampling algorithm to address the multi-agent Bayesian optimization problem under constrained communication. We develop a bound on Bayesian Simple Regret for this approach, where the bound is dependent on properties of the largest complete subgraph of the graph encoding communication structure between agents. With our bound, we show that in multi-agent communication networks, Bayesian Simple Regret will converge faster than in the sequential case (with the same number of rounds). Additionally, we demonstrate the efficacy of our algorithm with regret analysis on optimization test functions, illustrating faster convergence with graphs possessing larger connected subgraphs. Future work will focus on giving a tighter regret bound, and developing adjustments for the Distributed Thompson Sampling algorithm to better control the exploration/exploitation tradeoff.
Appendix A Additional Analysis
Lemma A.1.
([13][26]) At step $j,$ for all $x\in\mathcal{A},$ $\mathbb{E}[\mathbb{I}\{f(x)>U_{j}(x)\}\cdot(f(x)-U_{j}(x))]\leq\frac{1}{\sqrt{2\pi}}e^{-\beta_{j}/2}.$
Proof. Since $f$ is a $GP,$ we know $f(x)|D_{j}\sim\mathcal{N}(\mu_{j}(x),\sigma_{j}^{2}(x)).$ Recall that $U_{j}(\cdot)=\mu_{j}(\cdot)+\beta^{1/2}_{j}\sigma_{j}(\cdot).$ Thus, we know that $f(x)-U_{j}(x)|D_{j}\sim\mathcal{N}(-\beta^{1/2}_{j}\sigma_{j}(x),\sigma_{j}^{2}(x)).$ For a normal distribution $Z\sim\mathcal{N}(\mu,\sigma^{2}),$ with $\mu\leq 0,$ we have that $\mathbb{E}[Z\mathbb{I}(Z>0)]\leq\frac{\sigma}{\sqrt{2\pi}}e^{-\mu^{2}/(2\sigma^{2})}.$ Thus, we can apply this fact to our setting, yielding:
[…]
$\leq\frac{1}{\sqrt{2\pi}}e^{-\beta_{j}/2}$ $\text{(}\sigma_{j}(x)\leq\kappa(x,x)\leq 1\text{)}$
Lemma A.2.
[9] The information gain for selected points can be expressed in terms of the predictive variances. If $f_{[n]}=(f(x_{n}))\in\mathbb{R}^{n}:$
$I(y_{[n]};f_{[n]})=\frac{1}{2}\sum_{j=1}^{n}\log(1+\sigma_{\epsilon}^{-2}\sigma_{j-1}^{2}(x_{n}))$
Proof.
$I(y_{[n]};f_{[n]})$ $=H(y_{[n]})+H(y_{[n]}|f_{[n]})$ (Definition of information)
[…]
Now let’s develop an expression for $H(y_{[n]}):$
[…]
$=H(y_{[n-1]})+\frac{1}{2}\log(2\pi e(\sigma_{\epsilon}^{2}+\sigma^{2}_{n-1}(x_{n}))/2$ $\text{(}y_{n}|y_{[n-1]}=f(x_{n})+\epsilon_{n},y_{n}|y_{[n-1]}\sim N(\mu_{n-1}(x_{n}),\sigma^{2}_{n-1}(x_{n})+\sigma_{\epsilon}^{2}{)}$
We now have developed a recursive relation for $H(y_{[n]});$ we can inductively show that $H(y_{[n]})=\sum_{j=1}^{n}\frac{1}{2}\log(2\pi e(\sigma_{\epsilon}^{2}+\sigma^{2}_{j-1}(x_{n}))/2.$ Utilizing this expression, we can return to our mutual information expression:
$I(y_{[n]};f_{[n]})$ $=H(y_{[n]})-\frac{1}{2}\log|2\pi e\sigma_{\epsilon}^{2}\mathbf{I}|$ (Previously shown)
[…]
$=\frac{1}{2}\sum_{j=1}^{n}\log(1+\sigma_{j-1}^{2}(x_{n})\sigma_{\epsilon}^{-2})$ (Simplifying log subtraction)
Lemma A.3.
[13] Let $f\sim GP(0,\kappa),f:\mathcal{X}\to\mathbb{R}$ and each time we query $x\in\mathbb{X}$ we observe $y=f(x)+\varepsilon,$ where $\varepsilon\sim\mathcal{N}(0,\sigma_{n}^{2}).$ Let $\{x_{1},\ldots,x_{t}\}$ be an arbitrary set of $t$ evaluations to $f$ where $x_{j}\in\mathcal{X}$ for all $j.$ Let $\sigma_{j}^{2}$ denote the posterior variance conditioned on the first $j$ of these queries, $\{x_{1},\ldots,x_{j}\}.$ Then, $\sum_{j=1}^{n}\sigma_{j-1}^{2}(x_{j})\leq\frac{2\Psi_{n}}{\log(1+\sigma_{\epsilon}^{-2})}.$
Proof.
[…]
$\geq\frac{1}{2}\sum_{j=1}^{n}\log(1+\sigma_{\epsilon}^{-2}\sigma_{j-1}^{2}(x_{j}))$ (Lemma A.2)
Note that the function $\frac{x}{1+\log(x)}$ increases with $x.$ Also note that $0\leq\sigma_{\epsilon}^{2},\sigma_{j}^{2}(x_{j})\leq 1$ by assumption, and thus $\sigma_{\epsilon}^{-2}\sigma_{j}^{2}(x_{j})\leq\sigma_{\epsilon}^{-2}.$ Therefore,
$\frac{\sigma_{\epsilon}^{-2}\sigma_{j-1}^{2}(x_{j})}{\log(1+\sigma_{\epsilon}^{-2}\sigma_{j-1}^{2}(x_{j}))}\leq\frac{\sigma_{\epsilon}^{-2}}{\log(1+\sigma_{\epsilon}^{-2})}\implies\sigma_{j-1}^{2}(x_{j})\log(1+\sigma_{\epsilon}^{-2})\leq\log(1+\sigma_{\epsilon}^{-2}\sigma_{j-1}^{2}(x_{j})).$
Now we can revisit our earlier $\Psi_{n}$ expression:
$\Psi_{n}$ $\geq\frac{1}{2}\sum_{j=1}^{n}\log(1+\sigma_{\epsilon}^{-2}\sigma_{j-1}^{2}(x_{j}))$ (Lemma A.2)
$\geq\frac{1}{2}\log(1+\sigma_{\epsilon}^{-2})\sum_{j=1}^{n}\sigma_{j-1}^{2}(x_{j})$ (Previously shown)
With algebraic manipulation, the last expression is equivalent to $\sum_{j=1}^{n}\sigma_{j-1}^{2}(x_{j})\leq\frac{2\Psi_{n}}{\log(1+\sigma_{\epsilon}^{-2})},$ our desired statement.
Lemma A.4.
Let $f\sim GP(0,\kappa),$ and let $A,B$ be finite subsets of $\mathcal{X}.$ Let $y_{A}\in\mathbb{R}^{|A|}$ and $y_{B}\in\mathbb{R}^{|B|}$ denote the observations when we evaluate $f$ at $A$ and $B.$ Let $\sigma_{A},\sigma_{A\cup B}:\mathcal{X}\to\mathbb{R}$ denote the posterior standard deviation of the $GP$ when conditioned on $A$ and $A\cup B,$ respectively. Then
$\forall x\in\mathcal{X},\hskip 28.45274pt\frac{\sigma_{A}(x)}{\sigma_{A\cup B}(x)}=exp(I(f;y_{B}|y_{A}))$
Proof.
$I(f;y_{B}|y_{A})$ $=H(f|y_{A})-H(f|y_{A\cup B})$ (Mutual information with conditional entropy)
$=\frac{1}{2}\log|2\pi e\sigma_{A}^{2}|-\frac{1}{2}\log|2\pi e\sigma_{A\cup B}^{2}|$ (Entropy for GP)
$=\log\left(\frac{\sigma_{A}}{\sigma_{A\cup B}}\right)$
Thus, $\frac{\sigma_{A}}{\sigma_{A\cup B}}=exp(I(f;y_{B}|y_{A}))$
Appendix B Bounds for the information gain quantity $\Psi_{\tau}$ for different kernels
We note that following a known result in [9], $\Psi_{\tau}$ in fact satisfies sublinear growth for three well-known classes of kernels, namely the linear, exponential and Matern kernels.
Lemma B.1 (cf. Theorem 5 in [9]).
For any $\tau>0$, the maximal information gain $\Psi_{\tau}$ can be bounded as follows for the following kernels.
  1. 1.
    (Linear kernel): If $k(x,x^{\prime})=x^{\top}x^{\prime}$, then
    $\Psi_{\tau}=O(d\log(\tau)).$
  2. 2.
    (Squared exponential kernel): If $k(x,x^{\prime})=\exp(-\|x-x^{\prime}\|^{2}/2)$, then
    $\Psi_{\tau}=O(\left(\log(\tau)\right)^{d+1}).$
  3. 3.
    (Matern kernel with $\nu>1$): If $k(x,x^{\prime})=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{\sqrt{2\nu}}{d}\|x-x^{\prime}\|\right)^{v}K_{v}\left(\frac{\sqrt{2v}}{d}\|x-x^{\prime}\|\right)$, where $K_{v}(\cdot)$ is a modified Bessel function, and $\Gamma(\cdot)$ denotes the gamma function, then
    $\Psi_{\tau}=O((\tau)^{\frac{d(d+1)}{2\nu+d(d+1)}}\log(\tau))$
Appendix C Objective Functions
The test objective functions used in our simulations were the Rosenbrock and Ackley functions. The equations of those functions are as follows:
Rosenbrock:
$f(x,y)=(1-x)^{2}+100(y-x^{2})^{2}$ (19)
Ackley:
$f(x,y)=-20\exp\left(-0.2\sqrt{\frac{x^{2}+y^{2}}{2}}\right)-\exp\left(\frac{1}{2}(\cos(2\pi x)+\cos(2\pi y))\right)+20+\exp(1)$ (20)
(a) Plot of Rosenbrock Function.
(b) Plot of Ackley Function.
Figure 2: Plots of test objective functions.